The information provided below was taken from the books of Uli Enterprise for the financial year ended 31 May 2018. Extract of the statement of comprehensive income for the year ended 31 May 2018

$ Sales 480 000

Cost of sales 240 000

Rent income 12 000

Advertising 4 800

Salaries and wages 72 000

Rates and taxes 1 200

Other operating expenses 96 000

Additional information: 1. All inventories are priced at cost plus 25%.

2. Sales are divided equally each month. Sales are expected to increase by 18% for the financial year ending 31 May 2019.

3. Fifty percent (50%) of the sales are for cash and the balance is on credit. Debtors normally pay their accounts as follows:  40% in the month of the sale, and these debtors are entitled to a 5% discount;  55% one month after the sale The balance is usually written off as bad debts.

4. Purchases for June and July are expected to be 20 000 per month. All purchases are for cash.

5. In terms of the lease agreement rent is received monthly. The rent for the year ending 31 May 2019 is expected to be 10% more than the previous twelve months.

6. Advertising is paid for monthly and is estimated to be 8% of each month’s sales.

7. Salaries and wages will increase by 10% with effect from 01 July 2018. 8. Rates and taxes are paid in one installment for the year during July 2018. Rates are calculated at 80 cents (R0.80) per R100 on the value of the premises. The premises are valued at R300 000.

9. Other operating expenses are expected to increase by 5% and are spread evenly throughout the year. Operating expenses are paid for in the month in which they are incurred.

10. On 31 May 2018 an unfavorable bank balance of $18 000 appeared in the ledger of Uli Enterprise. Required: Prepare each of the following for June and July 2018 from the information provided above. (Provide separate monetary columns for each month.)

1.1 Debtors collection schedule

1.2 Cash budget
Answer #1
June July
Total Sale in 2018        40,000        40,000 (480000/12)
Total Sale in 2019        47,200        47,200 (40000*118%)
Cash sale-50%        23,600        23,600
Credit sale-50%        23,600        23,600
Credit sale-collection same month-5% discounted-40%           8,968           8,968 =23600*40%*95%
Credit sale-collection next month-55%        11,000        12,980 =40000*50%*55% =23600*55%
Debtor collection-Cash sale + Credit sale (same month)+credit sale(next month)        43,568        45,548
Cash budget
June July
Opening cash balance       (18,000)       (11,508)
Debtor collection        43,568        45,548
Purchases       (20,000)       (20,000)
Rental income (12000*110%/12)           1,100           1,100
Advertising-47200*8%         (3,776)         (3,776)
Salary & wages-72000/12, 72000/12*110%         (6,000)         (6,600)
Rates & Taxes- 300000*0.8/100                 -           (2,400)
Other operating expense-96000/12*105%         (8,400)         (8,400)
Cash closing balance      (11,508)        (6,036)